a firms C'=0.51q^2-7q+55.5. The values of its AFC at 20th unit of problem is 25. The demand function of this firm is 4q-200+4p=0. Should this firm opt for profit or sales maximization?
1
Expert's answer
2016-12-04T04:05:11-0500
TR=p*q=(50-q)*q=50q*q^2
MR=50-2q
MC=1.02q-7
50-2q=1.02q- 7
3.02q=57
Q=18.87